A stunt car travelling toward the edge of a cliff at 25 meters measured horizontally from the edge of the cliff.

A. How much time does it take the car to reach the ground?

B. Determine the height of the cliff.


1
Expert's answer
2021-05-27T12:29:27-0400

Gives

R=25m

time(t)=?

Hight(h)=?

Newton's 2nd law

"S=ut+\frac{1}{2}at^2"

u=0

a=g

"s=\frac{1}{2}gt^2"

"t=\sqrt\frac{2s}{g}"

Put value

"t=\sqrt\frac{2\times25}{9.8}" =2.259 sec

Newton's 3rd rules

"v^2=u^2+2as"

u=0

a=g

"v^2=2as"

"s=\frac{v^2}{2a}"

"v=\sqrt{2gs}"

"v=\sqrt{2\times9.8\times25}"

v=22.13m/sec

d"=u\times t"

d="22.13\times2.259=50m"